OpenStudy (anonymous):

I give medals, please! What is the solution to the equation? 1 1 1 - (3x-16) + -x = 5- 4 4 3 x=____

OpenStudy (anonymous):

@perl

OpenStudy (macielynn):

x=7 because -8 and 9 have like terms you can combine them -8+9=x x-21=-12-16 -12-16 is -28 x-21=-28 add -21 to both sides of the equation x=7

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

What about. 5(x+4) - 4x = 18 - 23 x=___

OpenStudy (anonymous):

@macielynn

OpenStudy (macielynn):

x=-2.7 repeating 5x+20-4x=18-23 add 4x to 5x 9x+20=18-23 18-23=-5 9x+20=-5 subtract 20 from both sides 9x=-25 you have to get 9 alone so you divide both sides by 9 -25 divided by 9 is -25/9 or you can reduce it to -2.7 repeating

OpenStudy (anonymous):

Thank you so much you cant believe how confusing this is for me and i might have 2 more if thats ok

OpenStudy (macielynn):

ok we learned this last year in 8th grade lol